# LoQT: Low Rank Adapters for Quantized Training

### v1

@@ Abstract
Training of large neural networks requires significant computational resources. Despite advances using low-rank adapters and quantization, pretraining of models such as LLMs on consumer hardware has not been possible without model sharding, offloading during training, or per-layer gradient updates. To address these limitations, we propose LoQT, a method for efficiently training quantized models. LoQT uses gradient-based tensor factorization to initialize low-rank trainable weight matrices that are periodically merged into quantized full-rank weight matrices. Our approach is suitable for both pretraining and fine-tuning models, which we demonstrate experimentally for language modeling and downstream task adaptation. We find that LoQT enables efficient training of models up to 7B parameters on a consumer-grade 24GB GPU. We also demonstrate the feasibility of training a 13B parameter model using per-layer gradient updates on the same hardware.
@@ 1 Introduction
Figure 1: Memory usage of Llama 13B, rank 1024. PL: per-layer gradient updates. A8bit: Adam 8-bit
Training large neural networks requires substantial hardware and energy resources. Reducing these requirements is thus important for cost efficiency and environmental sustainability, while also lowering the entry barrier for researchers and practitioners. The main barriers in training large models are the compute operations required, as well as the memory needed to store those computations, in this paper we focus on the latter. Memory use during training comes primarily from the weights of the model itself as well as the optimizer states used to train the model. To target the weights, variations on low-rank adaptation (LoRA) [16, 12, 6, 22, 23] have been suggested to decrease the number of trainable parameters, in combination with the use of low precision representations.To target the optimizer, low-rank approaches for projecting gradients to a lower rank have been used [44]. Finally, various applications of quantization [10, 26, 6] have been used to decrease memory requirements. In this work, we combine these approaches into a highly memory-efficient training configuration.
In typical training setups, the optimizer states take up larger space than the model itself, as methods such as Adam [18] need to keep track of two parameters for each weight of the model. GaLore [44] significantly reduces the number of parameters needed for storing the optimizer states by only keeping track of the optimizer state in a low-rank projection which is then projected up to be applied to the model weights. Combining this method with quantization would further shrink the footprint of the model. However, updating the weights of a highly quantized model directly in low-precision space has not been shown to work. This is mainly due to the higher-precision gradient updates being too small to have an impact on the lower-precision quantized states. Lastly, while LoRA is memory efficient for parameter-efficient fine-tuning of pre-trained models, it does not work as a pretraining method by itself [22].
@@
To address these shortcomings, we propose a new method, LoQT. LoQT initializes two low-rank factors for each weight matrix $W$: 1) $P$, using a projection of $W$’s gradients into a low-rank subspace, and 2) $B$, initialized to minimize the error of quantizing and represents the only low-rank matrix being trained. Only training $B$ means that the optimizer state can be shrunk significantly. The product $PB$ is periodically merged into the full rank matrix $W$ with exponentially increasing scheduling. As $W$ and $P$ do not receive gradient updates, they can be quantized, thus optimizing memory usage even further. We stress that it is the large accumulated updates that make it possible to update a quantized model as the addition of smaller changes would not register in the quantized state. A high-level overview is given in Fig. 2.
@@
We show that LoQT  works well with and without a quantized model, enabling not only a lower memory footprint in the optimizer state but also over the model parameters. Our results show that we get competitive performance to prior methods using significantly less memory, in particular when using quantization of the model weights in an application such as training an LLM. We also demonstrate superior performance when fine-tuning pre-trained models, by training and evaluating on the GLUE [36] benchmark for natural language understanding. Finally, we ablate several properties of the suggested approach and we find that an exponentially increasing projection gap is beneficial, not only to our work but also for GaLore. This is particularly crucial for the training of quantized models. LoQT  enables efficient training of 7B models on consumer-grade hardware with 24GB of memory, and makes it feasible to train models with up to 13 billion parameters without model parallel by making use of per-layer gradient updates [25] on the same hardware as shown in Fig. 1.
2 Related Work and Background
2.1 Neural Network Quantization and NF4
$w_{\text{NF4}}$ $=q_{\text{NF4}}(w,\mathcal{M}_{\beta})\coloneqq\text{argmin}_{q_{i}}|w/\mathcal{M}_{\beta}-q_{i}|,$ (1)
$w$ $=q_{\text{NF4}}^{-1}(w_{\text{NF4}},\mathcal{M}_{\beta})\coloneqq\mathcal{M}_{\beta}\cdot w_{\text{NF4}}.$ (2)
2.2 Adaptation of Pretrained Networks
Low-Rank Adaptation (LoRA) [16] enables fine-tuning of pre-trained models using low-rank adaptors, effectively reducing the memory footprint by only training weight adaptors for targeted layers. However, simple low-rank training using LoRA factor matrices has not been shown to work for pre-training[22].
2.3 Memory Efficient Optimization
Optimizer memory consumption
A significant portion of the memory needed to train neural networks is typically consumed by optimizer states. Notably, Adam [18], one of the most widely used optimizers, uses double the amount of memory as the gradient matrix to maintain first and second-order gradient statistics. Efforts to reduce this overhead have led to the development of adaptive optimization algorithms like Adafactor [32], which achieves sub-linear memory costs by factorizing the second-order statistics into a row-column outer product. GaLore [44] expands on this concept by using low-rank factorization and projecting low-rank gradients up to a full-rank when updating model weights.
Periodic updating of weight matrices
Low-rank gradients
@@ 3 Efficient Pretraining With LoQT
LoQT works by initializing and training low-rank adapters obtained by taking the SVD of a given layer’s gradients. Let $W$ indicate the full weights matrix of a given layer, $P$ be the left factor constructed from the SVD decomposition of the gradients matrix: $\nabla W=U\Sigma V^{\top}$; i.e. $P$ consists of the first $r$ columns of $U$ corresponding to the singular vectors with the $r$ largest singular values of $W$. The update rule for an interval $[T_{i-1},T_{i}]$ is then given by $W_{T_{i}}=W_{T_{i-1}}+PB$, where only the weights of $B$ are updated. $P$ and $W_{T_{i-1}}$ are kept constant over the time interval. We describe this in more detail below, followed by a discussion on periodic updating of the factor $P$, the enabling of quantized pre-training, error compensation, and exponential update intervals. Pseudo-code for LoQT is shown in Fig. 3.
@@ 3.1 Background: GaLore
Zhao et al. [44] show that gradients exhibit a low-rank structure during training. They exploit this insight by projecting the gradient to a low-rank subspace and applying the Adam optimizer before projecting back to the original dimensions. By doing this, the memory-intensive optimizer states required by Adam are shrunk significantly for low enough ranks.
@@
Gradient low-rank projection (GaLore) denotes the following gradient update rules, where $\eta$ is the learning rate, $\rho$ is the Adam optimizer, and $W\in R^{m\times n}$ is the weight matrix being trained, and $T$ represents the total number of training iterations between recomputing the projection matrices:
$W_{T}=W_{0}+\eta\sum_{t=0}^{T-1}\tilde{G}_{t},\text{ where}\hskip 10.00002pt\tilde{G}_{t}=P_{t}\rho_{t}(P_{t}^{\top}G_{t}Q_{t})Q_{t}^{\top},$
where $P_{t}\in R^{m\times r}$ and $Q_{t}\in R^{n\times r}$ are are the top-$r$ singular vectors from the SVD decomposition of the gradient matrix at each iteration $t$. In practice, this can be approximated by only applying a one-sided projection, as in
$W_{T}^{\prime}=W_{0}+\eta\sum_{t=0}^{T-1}P_{t}\rho_{t}(P_{t}^{\top}G_{t})\hskip 5.0pt\text{ or }\hskip 5.0ptW_{T}^{\prime}=W_{0}+\eta\sum_{t=0}^{T-1}\rho_{t}(G_{t}Q_{t})Q_{t}^{\top}.$
GaLore demonstrates that it is sufficient to keep the projection matrix fixed and only update it once every $T$ iterations, which we use in the following.
@@ 3.2 Low-rank Gradients as Adapters
We now describe the process by which we initialize the parameters we optimize in LoQT. We adopt the memory-performance trade-off of using only a one-sided projection. We compute $P^{\top}G$ if $m\leq n$ and $GQ$ otherwise. We want to achieve a separation between trainable weights and static weights, which we achieve by rewriting GaLore in terms of low-rank adaptors. Assume without loss of generality that $m\leq n$. Using the fact that $P_{t}$ is fixed in the interval $[0,T]$ we have that
@@
From (4) it is clear that we can keep track of low-rank updates using rank-$r$ low-rank adaptors. We note that in the interval $[0,T]$ only $B$ is updated, creating the desired separation. If implemented directly, we would need to compute the gradient with respect to $W$ and then project it down using $P^{\top}G_{t}$. We find that this step is unnecessary; it is sufficient to train $B$ using standard gradient descent.
We now show that training the $B$ matrix using gradient descent is equivalent to training w.r.t. $W_{t}$ as in definition 3.1 Let $G^{W}$ indicate the gradient of the loss with respect to $W$, and $G^{B}$ for the gradient of the loss with respect to $B$. Given a weight matrix $W$, an factor $P$ and a matrix $B$, when computing the forward pass $y=xW+xPB$, the gradient of a loss function $\mathcal{L}$ w.r.t. $B$ is $G^{B}=P^{\top}G^{W}$. This can be seen by applying the chain rule to get $G^{W}=x^{\top}\frac{\partial\mathcal{L}}{\partial y}$. The vector multiplied onto $B$ is $xP$ giving $G^{B}=(xP)^{\top}\frac{\partial\mathcal{L}}{\partial y}=P^{\top}x^{\top}\frac{\partial L}{\partial y}=P_{t}^{\top}G^{w}$. This shows that calculating the gradient w.r.t $B$ gives the same as projecting the gradient w.r.t $W$. It is thus clear that GaLore’s low-rank gradient updates should be the same as those obtained using backpropagation through LoRA.
3.3 Enabling pretraining with LoRA
Previous work has shown that training low-rank weight matrices works well for fine-tuning pre-trained weights. However, it has been shown that training low-rank factors, and periodically merging them into frozen $W$, does not work when starting with a randomly initialized matrix [22]. Here we address this shortcoming to enable full training using low-rank weight matrices.
@@ 3.3 Enabling pretraining with LoRA
$W_{t}=W_{0}+\Delta W_{T_{1}}+\Delta W_{T_{2}}+\ldots+\Delta W_{T_{n}},$
where $t=\sum_{i=1}^{|\mathcal{T}|}T_{i}$ and $\Delta W_{T_{i}}=P_{T_{i}}B_{T_{i}}$ represents the product of the learning from $B$ during the interval $T_{i}-T_{i-1}$ scaled by the learning rate $\eta$ and modulated by the gradient projection matrix $P_{T_{i}}$. After each update at iteration $T_{i}\in\mathcal{T}$, we reinitialize the low-rank factors $P_{T}$ and $B_{T}$. As in [44], we compute the gradient of $W_{T}$ over a single batch, focusing only on $\nabla W_{T}$ without needing full optimizer states. Not requiring optimizer states reduces the memory increase compared to full-rank training.
With the updated $W_{t}$ and reinitialized $P_{t}$ and $B_{t}$, a new gradient subspace is established for exploring the next $T_{i+1}-T_{i}$ steps. Our method treats $W_{t}$ as the full-rank repository of accumulated updates. Although it is periodically updated, $W_{t}$ is not part of the optimizer state computations and the gradients during the single forward pass are offloaded to cpu. Since the SVD calculations are done layerwise only the current layer is needed on GPU, or the SVD can be calculated on CPU. $P_{t}$ defines the general gradient subspace and trajectory for the upcoming $T_{i+1}-T_{i}$ steps, and $B_{t}$ is adjusted to navigate optimally within the direction set by $P_{t}$. As only $B_{t}$ is trained, the number of parameters needing optimizer states is drastically reduced.
@@ 3.4 Quantized Training
Given that $B$ is the only matrix accumulating gradients and undergoing changes, the other matrices $W$ and $P$ can be kept quantized. This approach allows storing the weights in NF4 precision without requiring high-precision gradient and weights to update $W$ and $P$. To the best of our knowledge, we are the first to enable efficient 4-bit quantized training using gradient descent without storing the weights in full precision.
We quantize weights $q_{\text{NF4}}(W)=W_{q}$ and $q_{\text{NF4}}(P)=P_{q}$ as described in §2.1. During periodic updates at interval time steps $(\sum_{i=1}^{n}T_{i})_{n=1}^{\text{max}}$, $P_{q}$ and $W_{q}$ are dequantized using the inverse function, $P_{\text{BF16}}=q_{\text{NF4}}^{-1}(P_{\text{NF4}})$ and $W_{BF16}=q_{\text{NF4}}^{-1}(W_{\text{NF4}})$. After this, $W_{T_{i}}=W_{T_{i-1}}+P_{T_{i-1}}B_{T_{i-1}}$ is computed and quantized. The quantization and dequantization processes are applied layer by layer, ensuring that not all layers are simultaneously in a non-quantized state to reduce memory usage. Moreover, the quantization state itself is re-quantized for further efficiency following [6]. We implement LoQT using weight-only quantization, meaning the quantized weights are loaded into memory and then dequantized before computing the matrix multiplications.
@@
0:  $W$: Weight, $T$: Update steps, $\eta$: LR, $r$: rank, $q_{N}(\cdot)$ and $deq_{N}(\cdot)$: N-bit quant and dequant functions.
@@
3.5 Compensating for Errors Introduced by Quantization
As the quantization process inevitably leads to a discrepancy between the non-quantized and quantized versions of $W$ we wish to reduce this effect as much as possible. While compensating for quantization errors has been done before [21], we need a tailored solution for LoQT.
During the merging phase, we first dequantize to obtain $W_{T-1}$ and $P_{T-1}$, and then compute the update $W_{T}=W_{T-1}+P_{T-1}B_{T-1}$. This is immediately followed by re-quantizing to get $Q_{T}=q_{\text{NF4}}(W_{T})$. Our goal is thus to minimize the quantization error $\|(Q_{T}+P_{T}B_{T})-W_{T}\|$. To achieve this, we solve for $B_{T}$ in the merging step, initializing $B_{T}$ as $B_{T}:=P_{T}^{+}(Q_{T}-W_{T})$, where $P_{T}^{+}$ is the Moore-Penrose pseudo-inverse. This approach avoids initializing $B_{T}$ as zeros and instead uses it for minimizing the quantization error $\|Q_{T}-W_{T}\|$. We then iteratively refine $B_{T}$, recomputing $Q_{T}=q_{\text{NF4}}(W_{T}-P_{T}B_{T})$, improving the alignment between the full-precision $W$ and its quantized state.
As training advances and the learning rate decays, the magnitude of the update $B_{T-1}$ to form $W_{T}$ decreases. This leads to negligible differences between $\left|q\left(Q_{t}+P_{t}B_{t}\right)-Q_{t}\right|$, which results in the weights plateauing early, as depicted in Fig. 4(a). To address this, we implement an exponentially increasing scheduler for updating $W$. Drawing from GaLore’s observation on the exponential decay of gradient rank (Lemma 3.1 [44]), we start with a frequency gap $\tau$ and progressively increase the update intervals by a factor of $\psi$. The sequence of updates is then given by $(T_{i})_{i=0}^{\infty}=(\tau+\psi^{i})_{i=0}^{\infty}$ Each $T_{i}$ marks a training step $t$ when $W$ is updated. This scheduling ensures more frequent updates earlier in training and more well-spaced adjustments later, allowing more accumulated gradients before each update.
4 Experiments and Results

### v4

# LoQT: Low-Rank Adapters for Quantized Pretraining

@@ Abstract
Despite advances using low-rank adapters and quantization, pretraining of large models on consumer hardware has not been possible without model sharding, offloading during training, or per-layer gradient updates. To address these limitations, we propose Low-Rank Adapters for Quantized Training (LoQT), a method for efficiently training quantized models. LoQT uses gradient-based tensor factorization to initialize low-rank trainable weight matrices that are periodically merged into quantized full-rank weight matrices. Our approach is suitable for both pretraining and fine-tuning models. We demonstrate this for language modeling and downstream task adaptation, finding that LoQT enables efficient training of models up to 7B parameters on a 24GB GPU. We also demonstrate the feasibility of training a 13B model using per-layer gradient updates on the same hardware.
@@ 1 Introduction
Figure 1: Memory usage of Llama 13B, rank 1024. LW: per-layer gradient updates. A8bit: Adam 8bit.
Training large neural networks requires substantial hardware and energy resources. Reducing these requirements is important for both cost efficiency and environmental reasons, while also lowering the entry barrier for researchers and practitioners in general. In this work, we target the memory component—a key part of the hardware requirements. Memory use during training comes primarily from storing the weights of the model, the optimizer states, and activations. To target the memory footprint of the weights, various applications of quantization [1, 2, 3, 4] have been used. For targeting the optimizer states, variations on low-rank adaptation (LoRA) [5, 6, 3, 7] have been suggested to decrease the number of trainable parameters for fine-tuning, in combination with the use of low precision representations. Low-rank approaches for projecting gradients to a lower rank have also been suggested [8]. In this work, we combine these approaches to address the model size and optimizer states, resulting in a highly memory-efficient configuration that is also suitable for pretraining.
@@
In typical training configurations, the optimizer states often take up more space than the model itself, as methods such as Adam [9] keep track of two parameters for each parameter of the model. While LoRA is memory efficient for parameter-efficient fine-tuning of pretrained models, it has not been shown to work as a pretraining method by itself [7]. GaLore [8] significantly reduces the memory needed for the optimizer parameters by storing the optimizer state in a low-rank projection, which is then projected up when applied to the model weights. Combining this method with quantization would further shrink the footprint of the model but this is not straightforward. Updating the weights of a highly quantized model directly in low-precision space has not been shown to work. This is mainly due to the higher-precision gradient updates having too small of an impact on the lower-precision quantized states.
@@
To address these shortcomings, we propose Low-Rank Adapters for Quantized Training (LoQT). LoQT initializes two low-rank factors, $P$ and $B$, for each weight matrix $W$. $P$ is initialized using a projection of $W$’s gradients into a low-rank subspace, and $B$ is initialized to minimize the quantization error. In our method, $B$ is the only matrix being actively optimized. Only optimizing $B$ means that the size of the gradients and optimizer state shrinks significantly compared to full training or LoRA. The product $PB$ is periodically merged into the full rank matrix $W$ with exponentially increasing gaps to account for smaller updates as the model converges, ensuring we accumulate large enough updates. As $W$ and $P$ do not receive gradient updates, they can be kept quantized, optimizing memory usage even further. It is the large accumulated updates that make it possible to update a quantized model—as the addition of smaller changes would not register in the quantized state. A high-level overview of our approach is given in Fig. 2.
@@
We show that LoQT works well both with and without quantization, enabling not only a lower memory footprint in the optimizer state but also over the model parameters. Our results show that we get competitive performance to prior methods using significantly less memory, in particular when quantizing the model weights in an application such as training a large language model (LLM). We also demonstrate comparable performance in language adaption, which we demonstrate on a curated Icelandic text dataset [10]. Finally, we show that LoQT also works for fine-tuning pretrained models on down-stream tasks, by training and evaluating on the GLUE [11] benchmark for natural language understanding and the GSM8K [12] dataset for mathematical reasoning. We ablate several properties of the suggested approach, demonstrating the importance of each component of LoQT. For instance, we find that an exponentially increasing projection gap is particularly crucial for the training of quantized models. An overview of memory savings is given in Fig. 1. We find that LoQT enables efficient training of 7B models on consumer-grade hardware with 24GB of memory, and makes it feasible to train models with up to 13 billion parameters without model parallelization, by making use of per-layer gradient updates [13].
@@ 2 Efficient Pretraining With LoQT
We now briefly introduce how LoQT works by initializing and training low-rank adapters. The adapters are initialized by taking the singular value decomposition (SVD) of a given layer’s gradients. We use $W$ to indicate the full weight matrix of a given layer and $P$ for the left factor constructed from the SVD decomposition of the gradient matrix, $\nabla W=U\Sigma V^{\top}$, such that $P$ consists of the first $r$ columns of $U$—corresponding to the singular vectors with the $r$ largest singular values of $W$, where $r$ is a given target rank. The update rule for a timestep $T_{i}$ is then given by $W_{T_{i}}=W_{T_{i-1}}+PB$. For the steps between $T_{i}$ and $T_{i+1}$ only the weights of $B$ are updated, while $P$ and $W_{T_{i-1}}$ remain constant. We describe this in more detail below, followed by a discussion on periodic updating of the factor $P$, enabling of quantized pretraining, error compensation, and exponential update intervals. Pseudo-code for LoQT is shown in Fig. 3.
@@ 2.1 Background: GaLore
Zhao et al. [8] find that gradients exhibit a low-rank structure during training. They exploit this insight by projecting the gradient to a low-rank subspace and applying the Adam optimizer before projecting back to the original dimensions. By doing this, the memory-intensive optimizer states required by Adam are shrunk significantly for low enough ranks.
@@
Gradient low-rank projection (GaLore) denotes the following gradient update rules, where $\eta$ is the learning rate, $\rho$ is the Adam optimizer, $W\in R^{m\times n}$ is the weight matrix being trained, and $T$ represents the total number of training iterations until the recomputation of the projection matrix:
$W_{T}=W_{0}+\eta\sum_{t=0}^{T-1}\tilde{G}_{t},\text{ where}\hskip 10.00002pt\tilde{G}_{t}=P_{t}\rho_{t}(P_{t}^{\top}G_{t}Q_{t})Q_{t}^{\top},$ (1)
where $r$ is a given target rank and $P_{t}\in R^{m\times r}$ and $Q_{t}\in R^{n\times r}$ are the top-$r$ singular vectors from the SVD decomposition of the gradient matrix at each iteration $t$. In practice, this can be approximated by only applying a one-sided projection, as in
$W_{T}^{\prime}=W_{0}+\eta\sum_{t=0}^{T-1}P_{t}\rho_{t}(P_{t}^{\top}G_{t})\hskip 5.0pt\text{ or }\hskip 5.0ptW_{T}^{\prime}=W_{0}+\eta\sum_{t=0}^{T-1}\rho_{t}(G_{t}Q_{t})Q_{t}^{\top}.$ (2)
Additionally, Zhao et al. [8] empirically show that it is sufficient to keep the projection matrix fixed and only update it once every $T$ iteration.
@@ 2.2 Low-rank Gradients as Adapters
We now describe how we initialize the parameters we optimize with LoQT. We start with the GaLore formulation from above and adopt the memory-performance trade-off of using only a one-sided projection (eq. 2), we compute $P^{\top}G$ if $m\leq n$ and $GQ$ otherwise. Our goal is to separate trainable weights and static weights, which we achieve by rewriting GaLore in terms of low-rank adaptors. We assume that $m\leq n$, if $m>n$ the same reasoning holds for $Q_{t}^{\top}$. Using the fact that $P_{t}$ is fixed on the interval $[0,T]$ we get
@@
It is clear from (4) that we can keep track of low-rank updates using rank-$r$ adaptors. We note that in the interval $[0,T]$ only $B$ is updated, creating the desired separation. If implemented directly, we would need to compute the gradient with respect to $W$ and then project it down using $P^{\top}G_{t}$. We find that this step is unnecessary; it is sufficient to train $B$ using standard gradient descent.
Equivalence of Gradient Updates
We point out that optimizing the low-rank matrix $B$ via gradient descent is equivalent to the projected gradient updates on $W_{t}$ described in Definition 2.1. Let $G^{W}=\frac{\partial\mathcal{L}}{\partial W}$ and $G^{B}=\frac{\partial\mathcal{L}}{\partial B}$ denote the loss gradients with respect to $W$ and $B$, respectively. Consider the forward pass $y=xW+xPB$, where $W$ is the weight matrix, $P$ is the projection matrix, and $B$ is the low-rank update matrix. By the chain rule:
$G^{B}$ $=(xP)^{\top}\frac{\partial\mathcal{L}}{\partial y}$ (5)
$=P^{\top}x^{\top}\frac{\partial\mathcal{L}}{\partial y}$ (6)
$=P^{\top}G^{W}$ (7)
This derivation establishes that computing gradients with respect to $B$ is equivalent to projecting the gradients with respect to $W$ onto the low-rank subspace defined by $P$. Therefore, GaLore’s low-rank gradient updates are identical to those obtained through backpropagation in LoRA.
2.3 Pretraining with LoRA
Previous work [5] has shown that training low-rank weight matrices works well for fine-tuning pretrained weights. However, it has been shown that starting with randomly initialized weights, training low-rank factors, and periodically merging them into a frozen weight matrix $W$, does not work when starting with a randomly initialized matrix [7]. We now address this to enable full training using low-rank weight matrices.
@@ 2.3 Pretraining with LoRA
$W_{t}=W_{0}+\Delta W_{T_{1}}+\Delta W_{T_{2}}+\ldots+\Delta W_{T_{n}},$ (8)
where $t=\sum_{i=1}^{|\mathcal{T}|}T_{i}$ and $\Delta W_{T_{i}}=P_{T_{i}}B_{T_{i}}$ represents the product of the learned matrix $B$ over the interval $T_{i}-T_{i-1}$ modulated by the gradient projection matrix $P_{T_{i}}$. After each periodic update at iterations $T_{i}\in\mathcal{T}$, we reinitialize the low-rank factors $P_{T}$ and $B_{T}$. As in [8], we compute the gradient of $W_{T}$ over a single batch, focusing only on $\nabla W_{T}$ without storing optimizer states for it, reducing the memory compared to full-rank training.
For each updated $W_{t}$ and reinitialized $P_{t}$ and $B_{t}$, a new gradient subspace is established for exploring the next $T_{i+1}-T_{i}$ steps. Our method treats $W_{t}$ as the full-rank repository of accumulated updates. Although it is periodically updated, $W_{t}$ is not part of the optimizer state computations, and the gradients during the single forward pass are offloaded to CPU/RAM. Since the SVD calculations are done layerwise, only the current layer needs to be on GPU, or the SVD can be calculated on CPU. $P_{t}$ defines the general gradient subspace and trajectory for the upcoming $T_{i+1}-T_{i}$ steps, and $B_{t}$ is adjusted to navigate within the direction set by $P_{t}$. As only $B_{t}$ is trained, the number of parameters requiring optimizer states is drastically reduced.
@@ 2.4 Quantized Training
Recall the update rule of our model, $W_{T_{i}}=W_{T_{i-1}}+PB$, given that $B$ is the only matrix accumulating gradients and undergoing changes, the other matrices $W$ and $P$ can be kept quantized. This approach allows storing the weights in NF4 precision [3] (see §5.1 for a detailed account) without requiring high-precision gradients and weights to update $W$ and $P$. To the best of our knowledge, we are the first to enable efficient 4-bit quantized pretraining using gradient descent without storing the weights in 16-bit precision.
We quantize the weights $q_{\text{NF4}}(W)=W_{q}$ and $q_{\text{NF4}}(P)=P_{q}$ as described in §5.1. During the periodic updates at interval time steps $(\sum_{i=1}^{n}T_{i})_{n=1}^{\text{max}}$, $P_{q}$ and $W_{q}$ are dequantized using the inverse function, $P_{\text{BF16}}=q_{\text{NF4}}^{-1}(P_{\text{NF4}})$ and $W_{BF16}=q_{\text{NF4}}^{-1}(W_{\text{NF4}})$. After this, $W_{T_{i}}=W_{T_{i-1}}+P_{T_{i-1}}B_{T_{i-1}}$ is computed and quantized. The quantization and dequantization processes are applied layer by layer, ensuring that not all layers are simultaneously in a non-quantized state to reduce memory usage. Moreover, the quantization state itself is re-quantized for further efficiency following [3]. We implement LoQT using weight-only quantization, this means that the quantized weights are loaded into memory and then dequantized before computing the matrix multiplications.
@@
0:  $W$: Weight, $T$: Update steps, $\eta$: LR, $r$: rank, $q_{N}(\cdot)$: N-bit quantization function.
@@
2.5 Compensating for Quantization Errors
As the quantization process inevitably results in rounding errors there is a discrepancy between the non-quantized and quantized versions of $W$. We wish to reduce this effect as much as possible. While compensating for quantization errors has been done before [14], we derive a tailored solution for LoQT.
[…]
As training advances and the learning rate decays, the magnitude of the update $B_{T-1}$ decreases. This leads to negligible differences $\left|q\left(Q_{t}+P_{t}B_{t}\right)-Q_{t}\right|$, which results in the loss plateauing early, as depicted in Fig. 4(a). To address this, we implement an exponentially increasing scheduler for updating $W$. Drawing from the observation that the gradient rank decays exponentially (Lemma 3.1 in  [8]), we start with an update interval $\tau$ and progressively increase the update intervals by a factor of $\psi$. The sequence of updates is then given by $(T_{i})_{i=0}^{\infty}=(\tau+\psi^{i})_{i=0}^{\infty}$. Each $T_{i}$ marks a training step $t$ when $W$ is updated. This scheduling ensures more frequent updates earlier in training and more well-spaced adjustments later, allowing for accumulation of sufficiently large gradients before each progressive update.
3 Experiments